How to find the equation of PX if the altitude through P(1,1) meets Q(5,4)R(4,0) in X?
Answer by rothauserc(4718)

the altitude is perpendicular to the line segment
:
slope of PX is the negative reciprocal of the line segment
:
slope of line segment = (0 - 4) / (4 - 5) = -4 / -1 = 4
:
slope of PX is -1/4
:
we are given (1, 1) is on PX
:
1 = (-1/4)1 + b, where b is the y-intercept
:
1 = -1/4 + b
:
b = 5/4
:
*******************************************
equation of PX in slope intercept form is
:
y = -x/4 + 5/4
:
equation of PX in standard form is
:
x + 4y - 5 = 0
******************************************
: